simplify (show your work)
-6^2 + (-1)^4

Found 2 solutions by Nate, fractalier:
Answer by Nate(3500)

-6^2 + (-1)^4
(-6)(-6) + (-1)(-1)(-1)(-1)
36 + 1
37

Answer by fractalier(6550)

-6^2 + (-1)^4 = -36 + 1 = -35
but if you mean
(-6)^2 + (-1)^4
then that equals
36 + 1 = 37